# Calibrated Language Models Must Hallucinate

### v1

@@ Abstract
Recent language models have a mysterious tendency to generate false but plausible-sounding text. Such “hallucinations” are an obstacle to the usability of language-based AI systems and can harm people who rely upon their outputs. This work shows shows that there is an inherent statistical reason that pretrained language models hallucinate certain types of facts, having nothing to do with the transformer LM architecture or data quality. For “arbitrary” facts whose veracity cannot be determined from the training data, we show that hallucination is necessary for language models that satisfy a statistical calibration condition appropriate for generative language models. Specifically, if the maximum probability of any fact is bounded, we show that the probability of generating a hallucination is close to the fraction of facts that occur exactly once in the training data (a “Good-Turing” estimate), even assuming ideal training data without errors.
One conclusion is that models pretrained to be sufficiently good predictors (i.e., calibrated) may require post-training to mitigate hallucinations on the type of arbitrary facts that tend to appear once in the training set. However, our analysis also suggests that there is no statistical reason that pretraining will lead to hallucination on facts that tend to appear more than once in the training data (like references to publications such as articles and books, whose hallucinations have been particularly notable and problematic) or on systematic facts (like arithmetic calculations). Therefore different architectures and learning algorithms may mitigate these latter types of hallucinations.
@@ 1 Introduction
The mysterious tendency of Language Models (LMs) to generate false information, such as references to non-existent article titles, has recently emerged as a critical issue. The popular term hallucination is defined in the Merriam-Webster (2023) dictionary as “a plausible but false or misleading response generated by an artificial intelligence algorithm.” In one case, lawyers were fined \$5,000 for submitting legal research containing hallucinated legal cases that they believed were correct (Shin, 2023). In healthcare, hallucinations could be life threatening to patients and physicians are concerned about malpractice cases (Mello and Guha, 2023). Furthermore, hallucinations have been widely reported on by the media (Weise and Metz, 2023), and the U.S. President recently put out an Executive Order calling for, among other things, safeguards against misleading outputs from generative AI systems (Biden, 2023). This paper presents a statistical justification for LM hallucination.
A language model is simply a probability distribution $D$ over sequences of tokens (e.g., words or other character sequences). Every distribution $D$ can equivalently be represented by its log-probabilities over entire sequences or conditional log-probabilities of the subsequent token given previous ones, $\log D(t_{1}\ldots t_{m})=\sum_{i=1}^{m}\log D(t_{i}\mid t_{1}\ldots t_{i-1}).$ This mathematical equivalence means that any LM can be used either to generate text or predict the next token in naturally occurring text conditioned on the previous tokens, with slightly different desiderata. For instance, consider the sentence,
@@
Sentences of this sort could be likely under a predictive LM, for example, to offer suggestions to reduce typing on phones (e.g., Tanaka-Ishii, 2007). It may be desirable to predict sandwich as an option of a word to type after the word tuna, along with other likely words such as salad and roll. On the other hand, the vast majority of sentences of this sort would be false if randomly fabricated by a generative LM. This paper shows LMs with good predictive text performance should hallucinate, even under ideal conditions. Notably in the first stage of “pretraining,” common today, the generative LM is optimized for predictive text performance (Chowdhery et al., 2022; OpenAI, 2023). Moreover, it gives a lower-bound on the rate of hallucination, which can shed light on the different rates at which different types of facts should be hallucinated.
What is common to both the potential references and the example above (which we shall refer to as 5W = Who-Ate-What-When-Where-Why factoids), above is that they are arbitrary in the sense that neither one can be determined systematically by rules—one cannot determine the veracity of most such facts that are not present in the training data. This in contrast to facts whose veracity can be determined systematically. We quantify how much LMs should hallucinate even in an simplified setting with several ideal properties. Because we are giving statistical lower-bounds, we favor simplicity over generality as the point of our lower bounds is to identify a fundamental cause of LM hallucination. Similar to classification, where one seeks a lower-bound for the difficulty of classification in noiseless settings (but noise-tolerant classifications algorithms), we seek a hallucination lower-bound that holds in the simplest setting where training data is i.i.d. without factual errors.
@@
Hallucinations have mystified LM users and researchers alike. Section 2 surveys numerous hypotheses that have been studied for why LMs hallucinate, ranging from inaccurate or outdated training data to the next-token log-likelihood objective in training. Hallucination can also be due to an adversarial or out-of-distribution prompt: a textual prefix provided for the LM to complete which establishes context. For example, there may be no factual completion to a fabricated prompt such as, The 15 Elves of San Salami are named…. [^2] A completion such as …actually never mind, I have no idea is unlikely given training data that does not have such retractions. In contrast, our work shows that even in an ideal world with perfect training data and no prompts, one should expect hallucinations from LMs which are calibrated.
@@ Results.
Our lower-bounds have the following interpretation. First, one should identify a large set of factoids: arbitrary, plausible, regular factoids. These could be posts about 5W and plausible research article citations. Intuitively, there are exponentially more plausible factoids that are incorrect (which we call hallucinations), than those that are facts. One can then consider what fraction of such factoids might appear exactly once in the training data. In the case of 5W, one could imagine that half of the posts occur exactly once. This would suggest that a calibrated-factoid model would hallucinate on about half of its generations on 5W factoids. On the other hand, one could imagine that there are many many fewer than $n$ articles and since the goal of publication is advertising, each reference might be expected to appear multiple times in the training data (i.e., have probability $\gg 1/n$) except for perhaps very recent publications (e.g., within the last day before they other references begin to appear). This would suggest that the missing mass for articles is low and that there is no inherent statistical necessity to hallucinate reference titles. There may be other causes of such hallucinations such as limited model capacity (even if the number of LM parameters is much greater than the number of articles, these parameters must encode many types of information beyond article titles). This also justifies the mitigation of consulting a fact database at generation time, even if that fact database is constructed solely from training data (Borgeaud et al., 2022).
@@
Following related work discussed in Section 2, Section 3 gives mathematical preliminaries including our definition of (mis)calibration. Section 4 defines our factoid model and Section 5 states the main results. Sections Section 6 proves the main lemma. Section 7 provides a simple upper bound showing that one cannot guarantee a hallucination rate greater than the monofact estimate for nearly calibrated estimators. Section 8 concludes and discusses limitations and future work. Appendix A in the appendix presents bounds on Good-Turing estimators from prior work. Appendix B discusses extensions to alternative models such as calibration using bins of equal width, those based on raw accuracy (cross-entropy), or those using using prompts. Appendix C gives additional proofs.
@@ 5 Guarantees
Fix any $\delta\in[0,1],b,n\in{\mathbb{N}},r,s\in{\mathbb{R}}$ and any $s$-sparse $D_{\mathrm{world}}$ with $r$-arbitrary-facts and $r$-arbitrary-probabilities. Then for any algorithm ${\mathcal{A}}:X^{n}\rightarrow\Delta(X)$, with probability $\geq 1-\delta$ over $D_{L}\sim D_{\mathrm{world}}$ and ${\bf x}_{\mathrm{train}}\sim D_{L}^{\times n}$,
@@
Analysis approach.
@@ 7 Upper bounds on hallucination rate
The algorithm is as follows.
@@ Appendix A Good-Turing estimator bounds
Eq. 13 is established by setting $\varepsilon=\sqrt{\ln(4/\delta)/n}$ in Eqs. 9 and 10, which by the union bound implies,
@@
Plugging $\delta^{\prime}=\delta/4$ in Eqs. 11 and 12 and the union bound gives,
@@
We first show Eq. 15. Note that Eq. 15 holds trivially for $n\leq 9$ because $GT(s),M_{\mathcal{D}}(s)\in[0,1]$ and $13\sqrt{\ln(4)/9}>$ 1. Thus, from Cor. 4, it suffices to verify that for $n>9,\delta\leq 1$:
@@
For Eq. 16, note that it holds trivially for $n\leq 6$ because $\sqrt{6\ln(2/\delta)/n}\geq\sqrt{6\ln(6)/6}>1$. Thus, from Cor. 4, it suffices to verify that for $n>6,\delta\leq 1/3$:
@@ B.3 Alternative calibration definitions
Appendix C Proofs of remaining corollaries
Rearranging terms,
It is now straightforward to prove the corollaries.
Proof of Corollary 1.
In the above we have used the fact that $O\subseteq F$ and $U\supseteq H$. The proof follows immediately from this, Eqs. 19 and 20 and the union bound, using $\Pi={\mathcal{V}}_{b}(g)$. ∎
==" alt="[LOGO]">

### v3

@@ Abstract
Recent language models generate false but plausible-sounding text with surprising frequency. Such “hallucinations” are an obstacle to the usability of language-based AI systems and can harm people who rely upon their outputs. This work shows that there is an inherent statistical lower-bound on the rate that pretrained language models hallucinate certain types of facts, having nothing to do with the transformer LM architecture or data quality. For “arbitrary” facts whose veracity cannot be determined from the training data, we show that hallucinations must occur at a certain rate for language models that satisfy a statistical calibration condition appropriate for generative language models. Specifically, if the maximum probability of any fact is bounded, we show that the probability of generating a hallucination is close to the fraction of facts that occur exactly once in the training data (a “Good-Turing” estimate), even assuming ideal training data without errors.
One conclusion is that models pretrained to be sufficiently good predictors (i.e., calibrated) may require post-training to mitigate hallucinations on the type of arbitrary facts that tend to appear once in the training set. However, our analysis also suggests that there is no statistical reason that pretraining will lead to hallucination on facts that tend to appear more than once in the training data (like references to publications such as articles and books, whose hallucinations have been particularly notable and problematic) or on systematic facts (like arithmetic calculations). Therefore, different architectures and learning algorithms may mitigate these latter types of hallucinations.
@@ 1 Introduction
The surprisingly high rate at which Language Models (LMs) generate false information, such as references to non-existent article titles, has recently emerged as a critical issue. The popular term hallucination is defined in the Merriam-Webster (2023) dictionary as “a plausible but false or misleading response generated by an artificial intelligence algorithm.” In one case, lawyers were fined \$5,000 for submitting legal research containing hallucinated legal cases that they believed were correct (Shin, 2023). In healthcare, hallucinations could be life threatening to patients and physicians are concerned about malpractice cases (Mello and Guha, 2023). Furthermore, hallucinations have been widely reported on by the media (Weise and Metz, 2023), and the U.S. President recently put out an Executive Order calling for, among other things, safeguards against misleading outputs from generative AI systems (Biden, 2023). This paper presents statistical lower-bounds on the rate of hallucination for LMs that are calibrated predictors of facts. This helps illuminate the nature of hallucination. It should not be taken to mean that hallucination is inevitable. Rather, as we discuss it is consistent with the fact that practitioners have increasingly been augmenting “pretraining” procedures by “post-training” procedures that reduce hallucination rates at the cost of reducing calibration as well.
An LM is simply a probability distribution $D$ over sequences of tokens, i.e., words or other character sequences. Clearly any LM which predicts every string with positive probability (a common property of LMs) will necessarily hallucinate with positive probability. However, if this probability is small, then hallucinations will be rare. Thus it is crucial to quantify the rate of hallucinations. Every distribution $D$ can equivalently be represented by its log-probabilities over entire sequences or conditional log-probabilities of the subsequent token given previous ones, $\log D(t_{1}\ldots t_{m})=\sum_{i=1}^{m}\log D(t_{i}\mid t_{1}\ldots t_{i-1}).$ This mathematically trivial equivalence [^1] The equivalence between next-token-prediction and document generation holds only if one disregards computational costs. It is similar to the statement that book can be output one word at a time, in order, in a single pass, even though writing a book generally involves many iterations. has a profound implication: any LM can be used either to generate text or predict the next token in naturally occurring text conditioned on the previous tokens, though prediction and generation have different desiderata. For instance, consider the sentence,
@@
Sentences of this sort could be likely under a predictive LM, for example, to offer suggestions to reduce typing on phones (e.g., Tanaka-Ishii, 2007). It may be desirable to predict sandwich as an option of a word to type after the word tuna, along with other likely words such as salad. On the other hand, the vast majority of sentences of this sort would be false if randomly fabricated by a generative LM. This paper shows LMs with good predictive text performance should hallucinate, even under ideal conditions. Notably in the first stage of pretraining, common today, the generative LM is optimized for predictive text performance (Chowdhery et al., 2022; OpenAI, 2023). Moreover, it gives a lower-bound on the rate of hallucination, which can shed light on the different rates at which different types of facts should be hallucinated.
What is common to both potential references and the example above (which we shall refer to as 5W = Who-Ate-What-When-Where-Why factoids), above is that they are arbitrary in the sense that neither one can be determined systematically by rules—one cannot determine the veracity of most such facts that are not present in the training data. This in contrast to facts whose veracity can be determined systematically. We quantify how much LMs should hallucinate even in an simplified setting with several ideal properties. Because we are giving statistical lower-bounds, we favor simplicity over generality as the point of our lower bounds is to identify a fundamental cause of LM hallucination. Similar to classification, where one seeks a lower-bound for the difficulty of classification in noiseless settings (but noise-tolerant classifications algorithms), we seek a hallucination lower-bound that holds in the simplest setting where training data is i.i.d. without factual errors.
@@
Hallucinations have mystified LM users and some researchers alike. Section 2 surveys numerous hypotheses that have been studied for why LMs hallucinate, ranging from inaccurate or outdated training data to the next-token log-likelihood objective in training. Hallucination can also be due to an adversarial or out-of-distribution prompt: a textual prefix provided for the LM to complete which establishes context. For example, there may be no factual completion to a fabricated prompt such as, The 15 Elves of San Salami are named…. [^3] A completion such as …actually never mind, I have no idea is unlikely given training data that does not have such retractions. In contrast, our work shows that even in an ideal, unchanging world with perfect training data and no prompts, one should expect hallucinations from LMs which are calibrated.
@@ Results.
Our lower-bounds have the following interpretation. First, one should identify a large set of factoids: arbitrary, plausible, regular factoids. These could be posts about 5W and plausible research article citations. Intuitively, there are exponentially more plausible factoids that are incorrect (which we call hallucinations), than those that are facts. One can then consider what fraction of such factoids might appear exactly once in the training data. In the case of 5W, one could imagine that half of the posts occur exactly once. This would suggest that a calibrated-factoid model would hallucinate on about half of its generations on 5W factoids. On the other hand, one could imagine that there are many many fewer than $n$ articles and since the goal of publication is advertising, each reference might be expected to appear multiple times in the training data [^4] Even papers that are never cited often appear on an authors’ CVs and institutional web pages and are cross-posted and listed in multiple publication indexes since the authors and institutions often wish to publicize their output and indices such as Google/Semantic Scholar may compete for completeness. (i.e., have probability $\gg 1/n$) except for perhaps very recent publications (e.g., within the last day before they other references begin to appear). This would suggest that the missing mass for articles is low and that there is no inherent statistical necessity to hallucinate reference titles. There may be other causes of such hallucinations such as limited model capacity (even if the number of LM parameters is much greater than the number of articles, these parameters must encode many types of information beyond article titles). This also justifies the mitigation of consulting a fact database at generation time, even if that fact database is constructed solely from training data (Borgeaud et al., 2022). Corollary 4 gives a simple generalization of our analysis to multiple types of facts.
Conversely, one may wonder what if any facts appear only once in a large training corpus that might encompass the entire web. First, it is worth noting that significant efforts are made, in constructing LM corpora, to remove duplicates (see, e.g. Shen et al., 2023). And in other contexts, such as Zipfian and other power-law distributions, it is often observed that a constant fraction of entities appear exactly once, even as the datasets scale. While most facts that come to mind may be well-known, many public meeting notes and other posts may contain unremarkable facts that are not mentioned in other places.
@@
Following related work discussed in Section 2, Section 3 gives mathematical preliminaries including our definition of (mis)calibration. Section 4 defines our factoid model and Section 5 states the main results. Sections Section 6 proves the main lemma. Section 7 provides a simple upper bound showing that one cannot guarantee a hallucination rate greater than the monofact estimate for nearly calibrated estimators. Section 9 concludes and discusses limitations and future work. Appendix A in the appendix presents bounds on Good-Turing estimators from prior work. Appendix B discusses extensions to alternative models such as calibration using bins of equal width, those based on raw accuracy (cross-entropy), or those using using prompts. Section 8 gives additional proofs.
@@ 5 Guarantees
Fix any $\delta\in[0,1],b,n\in{\mathbb{N}},r,s\in{\mathbb{R}}$ and any $s$-sparse $D_{\mathrm{world}}$ with $r$-regular-facts and $r$-regular-probabilities. Then for any algorithm ${\mathcal{A}}:X^{n}\rightarrow\Delta(X)$, with probability $\geq 1-\delta$ over $D_{L}\sim D_{\mathrm{world}}$ and ${\bf x}_{\mathrm{train}}\sim D_{L}^{\times n}$,
@@
5.1 Different types of facts
Our analysis immediately generalizes to multiple distinct types of facts (e.g., article references and social media posts). Suppose there are $k>1$ sets of factoids, $Y_{1},Y_{2},\ldots,Y_{k}$ and functions $f_{i}:X\rightarrow Y_{i}$, with corresponding sets of facts and hallucinations $F_{i}\cup H_{i}=Y_{i}$, monofact estimates $\widehat{MF}_{i}\in[0,1]$ and miscalibration rates $\mathrm{Mis}_{i,b}(g,p)$. One also would generalizes the notion of $s$-sparse to include the fact that, for each type of fact, $|F_{i}|\leq e^{-s}|H_{i}|$ with probability 1 and similarly generalize regular facts to hold for each type of fact. Then Corollary 1 implies:
Corollary 4.
Fix any $\delta\in[0,1],b,k,n\in{\mathbb{N}},s\in{\mathbb{R}}$ and any $s$-sparse regular $D_{\mathrm{world}}$. Then for any algorithm ${\mathcal{A}}:X^{n}\rightarrow\Delta(X)$, with probability $\geq 1-\delta$ over $D_{L}\sim D_{\mathrm{world}}$ and ${\bf x}_{\mathrm{train}}\sim D_{L}^{\times n}$,
$g(H_{i})\geq\widehat{MF}_{i}-\mathrm{Mis}_{i,b}(g,p)-\frac{3ke^{-s}}{\delta}-\sqrt{\frac{6\ln(6k/\delta)}{n}}\text{ for }i=1,2,\ldots,k.$
The proof follows trivially from Corollary 1 using the union bound and $\delta/k$ failure probability for each type of fact.
5.2 Analysis approach
@@ 7 Upper bounds on hallucination rate
The algorithm memorizes the facts in the training data and generates random facts from the training data and random unseen factoids uniformly at random, but at different rates.
@@
8 Proofs of Corollaries
Rearranging terms,
It is now straightforward to prove the corollaries.
Proof of Corollary 1.
In the above we have used the fact that $O\subseteq F$ and $U\supseteq H$. The proof follows immediately from this, Eqs. 9 and 10 and the union bound, using $\Pi={\mathcal{V}}_{b}(g)$. ∎
@@ 9 Conclusions, limitations, and future work
Acknowledgments.
We thank Mirac Suzgun and Kevin Leyton-Brown for helpful discussions.
@@ Appendix A Good-Turing estimator bounds
Eq. 15 is established by setting $\varepsilon=\sqrt{\ln(4/\delta)/n}$ in Eqs. 11 and 12, which by the union bound implies,
@@
Plugging $\delta^{\prime}=\delta/4$ in Eqs. 13 and 14 and the union bound gives,
@@
We first show Eq. 17. Note that Eq. 17 holds trivially for $n\leq 9$ because $GT(s),M_{\mathcal{D}}(s)\in[0,1]$ and $13\sqrt{\ln(4)/9}>$ 1. Thus, from Cor. 5, it suffices to verify that for $n>9,\delta\leq 1$:
@@
For Eq. 18, note that it holds trivially for $n\leq 6$ because $\sqrt{6\ln(2/\delta)/n}\geq\sqrt{6\ln(6)/6}>1$. Thus, from Cor. 5, it suffices to verify that for $n>6,\delta\leq 1/3$:
@@ B.2 KL-divergence and log-loss
Nonetheless, it would be interesting to see if one can quantify hallucination rates purely on the basis of accuracy rather than calibration. A starting point may be the work of Feldman (2021) which shows that statistical models must memorize their training data for accuracy purposes. [^5] We acknowledge Gautam Kamath for pointing out this connection.